# Imaging Shapes of Atomic Nuclei in High-Energy Nuclear Collisions

### v1

@@ Methods
During the offline analysis, events are selected to have collision vertices $z_{\mathrm{vtx}}$ within 30 cm of the TPC center along the beam line, and within 2 cm of the beam spot in the transverse plane. Additionally, a selection criterion based on the correlation between the number of TPC tracks and the number of tracks matched to the time-of-flight detector covering $|\eta|<0.9$ is applied to suppress pileup events (events containing more than one collision in the TPC) Llope (2012) and background events.
After applying these selection criteria, the Au+Au data set consists of approximately 528 million minimum-bias events (including 370 million from 2011) and 120 million UCC events. The U+U data set comprises around 300 million minimum-bias events and 5 million UCC events.
.4 Track selection
The charged particle tracks are reconstructed within $|\eta|<1$ using the TPC Anderson et al. (2003). For this analysis, tracks are selected with $|\eta|<1$ and the transverse momentum range $3.00.2<p_{\mathrm{T}}<$ 3.0 GeV/$c$. To ensure good quality, the selected tracks are required to have at least 16 fit points out of a maximum of 45, and the ratio of the number of fit points to the number of possible points must be greater than 0.52. Additionally, to reduce contributions from secondary decays, the distance of the track’s closest approach (DCA) to the primary collision vertex is required to be less than 3 cm.
The tracking efficiency in the TPC was evaluated using the standard STAR Monte Carlo embedding technique Adamczyk et al. (2015b). The efficiencies are nearly independent of $p_{\mathrm{T}}$ for $p_{\mathrm{T}}>0.5$ GeV/$c$, with plateau values ranging from 0.72 (0.69) in the most central Au+Au (U+U) collisions to 0.92 in the most peripheral collisions. The efficiency exhibits some $p_{\mathrm{T}}$-dependent variation, on the order of 10% of the plateau values, within the range of $0.50.2<p_{\mathrm{T}}<$ 0.5 GeV/$c$.
.5 Centrality
The centrality of each collision is determined using $N_{\mathrm{ch}}^{\mathrm{rec}}$, which represents the number of raw reconstructed tracks in $|\eta|<0.5$, satisfying $p_{\mathrm{T}}>0.15$ GeV/$c$, and having more than 10 fit points. After applying a correction to account for the dependence on the collision vertex position and the luminosity, the distribution of $N_{\mathrm{ch}}^{\mathrm{rec}}$ is compared to a Monte Carlo Glauber calculation Adamczyk et al. (2015b). This comparison allows for the determination of centrality intervals expressed as a percentage of the total nucleus-nucleus inelastic cross-section.
.6 Calculation of observables
@@
$\left\langle v_{2}^{2}\right\rangle$ $=\left\langle\frac{\sum_{i\neq j}w_{i}w_{j}\cos 2(\phi_{i}-\phi_{j})}{\sum_{i\neq j}w_{i}w_{j}}\right\rangle_{\mathrm{evt}}$
@@
The averages are performed first on all multiplets within a single event and then over all events in a fixed $N_{\mathrm{ch}}^{\mathrm{rec}}$ bin. The track-wise weights $w_{i,j,k}$ account for tracking efficiency and its $\eta$ and $\phi$-dependent variation. The values of $\left\langle v_{2}^{2}\right\rangle$ and $\left\langle(\delta p_{\mathrm{T}})^{2}\right\rangle$ are obtained using the standard method, where particle $i$ and $j$ are selected from the $|\eta|<1$, as well as the two-subevent method, where particle $i$ and $j$ are selected from pseudorapidity ranges of $-1<\eta_{i}<-0.1$ and $10.1<\eta_{j}<$ 1, respectively.
@@
The covariance $\left\langle v_{2}^{2}\delta p_{\mathrm{T}}\right\rangle$ is calculated by averaging over all triplets labeled by particle indices $i$, $j$, and $k$. The standard cumulant framework is used to obtain the results instead of directly calculating all triplets Bilandzic et al. (2011). We also calculated $\left\langle v_{2}^{2}\delta p_{\mathrm{T}}\right\rangle$ using the two- and three-subevent methods Aad et al. (2023). In the two-subevent method, particles $i$ and $j$ are taken from ranges of $-1<\eta_{i}<-0.1$ and $10.1<\eta_{j}<$ 1, while particle $k$ can be taken from either subevents. In the three-subevent method, all three particles $i$, $j$ and $k$ are taken from distinct $\eta$ ranges, namely $-1<\eta_{i}<-0.4$, $10.4<\eta_{j}<$ 1, and $|\eta_{k}|<0.3$. The inclusion of a pseudorapidity gap between the particle pairs or triplets suppresses the short-range “non-flow” correlations arising from resonance decays and jets Jia et al. (2017).
@@
The calculation of $\rho_{2}~{}=\frac{\left\langle v_{2}^{2}\delta p_{\mathrm{T}}\right\rangle}{\left\langle v_{2}^{2}\right\rangle\sqrt{\left\langle(\delta p_{\mathrm{T}})^{2}\right\rangle}}$ relies on the input values of $\left\langle v_{2}^{2}\right\rangle$, $\left\langle(\delta p_{\mathrm{T}})^{2}\right\rangle$ and $\left\langle v_{2}^{2}\delta p_{\mathrm{T}}\right\rangle$. These components and $\rho_{2}$ are shown in Fig. 4 as a function of centrality. In the central region, enhancements of $\left\langle v_{2}^{2}\right\rangle$ and $\left\langle(\delta p_{\mathrm{T}})^{2}\right\rangle$ are observed in U+U collisions relative to Au+Au collisions, which is consistent with the influence of large $\beta_{2\rm U}$. In contrast, the values of $\left\langle v_{2}^{2}\delta p_{\mathrm{T}}\right\rangle$ are significantly suppressed in U+U compared to Au+Au collisions across the entire centrality range shown. This suppression is consistent with the negative contribution expected for strong prolate deformation of U as described in Eq. (2).
.7 Influence of non-flow correlations
The results in Fig. 4 are obtained using the two-subevent method. However, as discussed earlier, $\left\langle v_{2}^{2}\delta p_{\mathrm{T}}\right\rangle$ has contributions from non-flow correlations, which can be explored by comparing results obtained from the standard, two-subevent and three-subevent methods. The results for $\rho_{2}$ are shown in Fig. 5. In this figure, $\left\langle v_{2}^{2}\right\rangle$ and $\left\langle(\delta p_{\mathrm{T}})^{2}\right\rangle$ are calculated using the two-subevent method, so any variations in $\rho_{2}$ arise only from $\left\langle v_{2}^{2}\delta p_{\mathrm{T}}\right\rangle$.
Differences are observed between the standard and subevent methods, particularly in peripheral collisions. The majority of the variations arise from the comparison between the standard and two-subevent methods, while the differences between the two- and three-subevent methods are relatively small. It is worth mentioning that some of these differences can be attributed to longitudinal fluctuations in $[p_{\mathrm{T}}]$ and $v_{2}$ Khachatryan et al. (2015); Aad et al. (2021, 2023).
.8 Systematic uncertainties
@@
Systematic uncertainties are estimated by varying the track quality selections, the $z_{\mathrm{vtx}}$ cuts, examining the influence of pileup, comparing results from periods with different detector conditions, the possible impact of non-flow contributions, and closure test. The influence of track selection criteria is studied by varying the number of fit hits on the track from a minimum of 16 to 19 and by varying DCA cut from $<3$ cm to $<2.5$ cm, resulting in variations of 1–5% for $\left\langle(\delta p_{\mathrm{T}})^{2}\right\rangle$. The impacts on $\left\langle v_{2}^{2}\right\rangle$ and $\left\langle v_{2}^{2}\delta p_{\mathrm{T}}\right\rangle$ are up to $2.5\%$ and $4\%$, respectively.
@@
The influence of track reconstruction on the collision vertex is examined by comparing the results for different $|z_{\mathrm{vtx}}|$ cuts, with differences found to be $0.5$–3% for all observables. Comparisons between different data-taking periods, particularly normal and reverse magnetic field runs in Au+Au collisions, show consistency within their statistical uncertainties. The influence of pileup and background events is studied by varying the cut on the correlation between $N_{\mathrm{ch}}^{\mathrm{rec}}$ and the number of hits in the TOF. The influence is found to be 1–3% for $\left\langle v_{2}^{2}\right\rangle$ and $\left\langle(\delta p_{\mathrm{T}})^{2}\right\rangle$, and reaches 2–10% for $\left\langle v_{2}^{2}\delta p_{\mathrm{T}}\right\rangle$. Comparisons are also made between the 2010 and 2011 Au+Au datasets, which have different active acceptances in the TPC. The results are largely consistent with the quoted uncertainties, although some differences are observed, particularly in the central region, where variations reach 5–10% for $\left\langle v_{2}^{2}\delta p_{\mathrm{T}}\right\rangle$.
@@
A closure test was conducted, wherein the reconstruction efficiency and its variations in $\eta$ and $\phi$ from the data were utilized to retain a fraction of the particles generated from a multi-phase transport model Lin et al. (2005). Subsequently, a track-by-track weight, as described in Eq. 9, was applied to the accepted particles. All observables are calculated using the accepted particles and compared to those obtained using the original particles. This procedure allowed us to recover $\left\langle v_{2}^{2}\right\rangle$ and $\left\langle(\delta p_{\mathrm{T}})^{2}\right\rangle$ within their statistical uncertainties. However, a 2–3% nonclosure was observed in $\left\langle v_{2}^{2}\delta p_{\mathrm{T}}\right\rangle$. Nevertheless, it’s important to note that such non-closures largely cancel out when considering the ratios between U+U and Au+Au collisions.
@@
Several additional cross-checks were also carried out. The track reconstruction efficiency has an approximate 5% uncertainty due to its reliance on particle type and occupancy dependence. We repeated the analysis by varying this efficiency, and the variations in the results were either less than 1% or consistent with their statistical uncertainties. The reconstructed $p_{\mathrm{T}}$ can differ from the true value due to finite momentum resolution. This effect was investigated by further smearing the reconstructed $p_{\mathrm{T}}$ according to the known resolution, calculating the observable, and comparing the results with the original ones. A discrepancy of approximately 0.5% was observed for $\left\langle(\delta p_{\mathrm{T}})^{2}\right\rangle$, while other observables remained consistent within their statistical uncertainties. These effects cancel in the ratios between U+U and Au+Au collisions.
Modest differences are observed between the standard method and subvent methods for all observables, particularly towards the peripheral collisions. Some of these differences are attributable to non-flow contributions, while the remainder could arise from longitudinal flow decorrelations. We obtain default results as the average of the standard method and two-subevent method, and half of the differences are assigned as systematic uncertainties. They are 1–2% for $\left\langle v_{2}^{2}\right\rangle$, 1–3% for $\left\langle(\delta p_{\mathrm{T}})^{2}\right\rangle$, and 2–4% for $\left\langle v_{2}^{2}\delta p_{\mathrm{T}}\right\rangle$ in the 0–40% centrality range, respectively.

### v2

@@ Methods
In the offline analysis, events are selected to have collision vertices $z_{\mathrm{vtx}}$ within 30 cm of the TPC center along the beamline and within 2 cm of the beam spot in the transverse plane. Additionally, a selection criterion based on the correlation between the number of TPC tracks and the number of tracks matched to the time-of-flight detector covering $|\eta|<0.9$ is applied to suppress pileup events (events containing more than one collision in the TPC) Llope (2012) and background events.
After applying these selection criteria, the Au+Au dataset has approximately 528 million minimum-bias events (including 370 million in 2011) and 120 million UCC events. The U+U dataset comprises around 300 million minimum-bias events and 5 million UCC events.
.5 Track selection
For this analysis, tracks are selected with $|\eta|<1$ and the transverse momentum range $3.00.2<p_{\mathrm{T}}<$ 3.0 GeV/$c$. To ensure good quality, the selected tracks must have at least 16 fit points out of a maximum of 45, and the ratio of the number of fit points to the number of possible points must be greater than 0.52. Additionally, to reduce contributions from secondary decays, the distance of the track’s closest approach (DCA) to the primary collision vertex must be less than 3 cm.
The tracking efficiency in the TPC was evaluated using the standard STAR Monte Carlo embedding technique Adamczyk et al. (2015). The efficiencies are nearly independent of $p_{\mathrm{T}}$ for $p_{\mathrm{T}}>0.5$ GeV/$c$, with plateau values ranging from 0.72 (0.69) in the most central Au+Au (U+U) collisions to 0.92 in the most peripheral collisions. The efficiency exhibits some $p_{\mathrm{T}}$-dependent variation, on the order of 10% of the plateau values, within the range of $0.50.2<p_{\mathrm{T}}<$ 0.5 GeV/$c$.
.6 Centrality
The centrality of each collision is determined using $N_{\mathrm{ch}}^{\mathrm{rec}}$, which represents the number of raw reconstructed tracks in $|\eta|<0.5$, satisfying $p_{\mathrm{T}}>0.15$ GeV/$c$, and having more than 10 fit points. After applying a correction to account for the dependence on the collision vertex position and the luminosity, the distribution of $N_{\mathrm{ch}}^{\mathrm{rec}}$ is compared with a Monte Carlo Glauber calculation Adamczyk et al. (2015). This comparison allows for determining centrality intervals, expressed as a percentage of the total nucleus-nucleus inelastic cross-section.
.7 Calculation of observables
@@
$\left\langle v_{2}^{2}\right\rangle$ $=\left\langle\frac{\sum_{i\neq j}w_{i}w_{j}\cos(2(\phi_{i}-\phi_{j}))}{\sum_{i\neq j}w_{i}w_{j}}\right\rangle_{\mathrm{evt}}$
@@
The averages are performed first on all multiplets within a single event and then over all events in a fixed $N_{\mathrm{ch}}^{\mathrm{rec}}$ bin. The track-wise weights $w_{i,j,k}$ account for tracking efficiency and its $\eta$ and $\phi$ dependent variations. The values of $\left\langle v_{2}^{2}\right\rangle$ and $\left\langle(\delta p_{\mathrm{T}})^{2}\right\rangle$ are obtained using the standard method, in which particle $i$ and $j$ are selected from $|\eta|<1$, as well as the two-subevent method, in which particle $i$ and $j$ are selected from pseudorapidity ranges of $-1<\eta_{i}<-0.1$ and $10.1<\eta_{j}<$ 1, respectively. We also calculate the efficiency-corrected charged particle multiplicity in $|\eta|<0.5$, defined as $N_{\mathrm{ch}}=\sum_{i}w_{i}$. This observable is used for evaluating the systematics.
@@
The covariance $\left\langle v_{2}^{2}\delta p_{\mathrm{T}}\right\rangle$ is calculated by averaging over all triplets labeled by particle indices $i$, $j$, and $k$. The standard cumulant framework is used to obtain the results instead of directly calculating all triplets Bilandzic et al. (2011). We also calculated $\left\langle v_{2}^{2}\delta p_{\mathrm{T}}\right\rangle$ using the two-subevent method Aad et al. (2023), where particles $i$ and $j$ are taken from ranges of $-1<\eta_{i}<-0.1$ and $10.1<\eta_{j}<$ 1, while particle $k$ is taken from either subevents. Including a pseudorapidity gap between the particle pairs or triplets suppresses the short-range “non-flow” correlations arising from resonance decays and jets Jia et al. (2017).
@@
The calculation of $\rho_{2}~{}=\frac{\left\langle v_{2}^{2}\delta p_{\mathrm{T}}\right\rangle}{\left\langle v_{2}^{2}\right\rangle\sqrt{\left\langle(\delta p_{\mathrm{T}})^{2}\right\rangle}}$ relies on the input values of $\left\langle v_{2}^{2}\right\rangle$, $\left\langle(\delta p_{\mathrm{T}})^{2}\right\rangle$ and $\left\langle v_{2}^{2}\delta p_{\mathrm{T}}\right\rangle$. These components and $\rho_{2}$ are shown in Fig. 4 as a function of centrality. In the central region, enhancements of $\left\langle v_{2}^{2}\right\rangle$ and $\left\langle(\delta p_{\mathrm{T}})^{2}\right\rangle$ are observed in U+U relative to Au+Au collisions, which is consistent with the influence of large $\beta_{2\rm U}$. In contrast, the values of $\left\langle v_{2}^{2}\delta p_{\mathrm{T}}\right\rangle$ are significantly suppressed in U+U compared with Au+Au collisions across the entire centrality range shown. This suppression is consistent with the negative contribution expected for strong prolate deformation of U as described in Eq. (2).
Fig. 4: Components involved in $v_{2}$-$[p_{\mathrm{T}}]$ correlations. The centrality dependences of $\left\langle v_{2}^{2}\right\rangle$ (a), $\left\langle(\delta p_{\mathrm{T}})^{2}\right\rangle$ (b), $\left\langle v_{2}^{2}\delta p_{\mathrm{T}}\right\rangle$ (c) and $\rho_{2}~{}=~{}\frac{\left\langle v_{2}^{2}\delta p_{\mathrm{T}}\right\rangle}{\left\langle v_{2}^{2}\right\rangle\sqrt{\left\langle(\delta p_{\mathrm{T}})^{2}\right\rangle}}$ (d) in U+U (red) and Au+Au (black) collisions. They are calculated using the two-subevent method. The error bars and shaded bands are statistical and systematic uncertainties, respectively.
In this analysis, the default results are obtained using the two-subevent method. The differences between the standard and two-subevent methods are used to evaluate the impact of non-flow correlations discussed below.
.8 Influence of non-flow correlations
An important background in our measurement is non-flow: correlations among a few particles originated from a common source, such as resonance decays and jets, which are uncorrelated with the initial geometry. Two approaches are used to estimate the non-flow contributions. Non-flow correlations are short-range in $\eta$ and can be suppressed via the subevent method by requiring a rapidity gap between the pairs or triplets of particles in Eq.(9). Hence, in the first approach, the differences between the standard and subevent methods provide an estimate of the non-flow contribution. However, one must keep in mind that part of the rapidity gap dependence of the signal in central collisions may arise from longitudinal fluctuations in $[p_{\mathrm{T}}]$ and $v_{2}$ due to variations in the initial geometry in $\eta$ Aad et al. (2023).
The second approach assumes that the clusters causing non-flow correlations are mutually independent. In this independent-source scenario, non-flow in $n$-particle cumulants is expected to be diluted by the charged particle multiplicity as $1/N_{\mathrm{ch}}^{n-1}$ Borghini et al. (2001). Therefore, non-flow (nf) contributions in a given centrality can be estimated by,
$\left\langle v_{2}^{2}\right\rangle_{\mathrm{nf}}\approx\frac{\left[\left\langle v_{2}^{2}\right\rangle N_{\mathrm{ch}}\right]_{\mathrm{peri}}}{N_{\mathrm{ch}}}\;,$
$\left\langle(\delta p_{\mathrm{T}})^{2}\right\rangle_{\mathrm{nf}}\approx\frac{\left[\left\langle(\delta p_{\mathrm{T}})^{2}\right\rangle N_{\mathrm{ch}}\right]_{\mathrm{peri}}}{N_{\mathrm{ch}}}\;,$
$\left\langle v_{2}^{2}\delta p_{\mathrm{T}}\right\rangle_{\mathrm{nf}}\approx\frac{\left[\left\langle v_{2}^{2}\delta p_{\mathrm{T}}\right\rangle N_{\mathrm{ch}}^{2}\right]_{\mathrm{peri}}}{N_{\mathrm{ch}}^{2}}$ (10)
where the subscript “peri” is a label for the peripheral bin. This procedure makes two assumptions that are not fully valid: 1) the signal in the peripheral bin is all non-flow, and 2) non-flow in other centralities is unmodified by final state medium effects. For example, the medium effects strongly suppress the jet yield and modify the azimuthal structure of non-flow correlations. Hence, this approach only provides a qualitative estimate of the non-flow. Moreover, this approach is not applicable for $\left\langle(\delta p_{\mathrm{T}})^{2}\right\rangle$, as medium effects are expected to reduce the momentum differences of non-flow particles, since they are out of local equilibrium.
Figure 5 shows the $N_{\mathrm{ch}}$-scaled values of $\left\langle v_{2}^{2}\right\rangle$, $\left\langle(\delta p_{\mathrm{T}})^{2}\right\rangle$, and $\left\langle v_{2}^{2}\delta p_{\mathrm{T}}\right\rangle$ as a function of centrality in Au+Au collisions. The requirement of subevent reduces the signal in the most peripheral bin by 50%, 40%, and 80%, respectively, which can be treated as the amount of non-flow rejected by the subevent requirement. Therefore, we use the differences between the standard and subevent methods to estimate the non-flow in the subevent method. These differences vary with centrality because of the combined effects of medium modification of non-flow and longitudinal flow decorrelations Aad et al. (2021). These differences are propagated to the ratios of these observables between U+U and Au+Au. They are found to be 1.1%, 3.5%, and 11% for $R_{v_{2}^{2}}$, $R_{(\delta p_{\mathrm{T}})^{2}}$, and $R_{v_{2}^{2}\delta p_{\mathrm{T}}}$, respectively.
Fig. 5: Impact of non-flow correlations. The centrality dependence of $N_{\mathrm{ch}}$ scaled quantities in Au+Au collisions, $\left\langle v_{2}^{2}\right\rangle N_{\mathrm{ch}}$ (a), $\left\langle(\delta p_{\mathrm{T}})^{2}\right\rangle N_{\mathrm{ch}}$ (b), and $\left\langle v_{2}^{2}\delta p_{\mathrm{T}}\right\rangle N_{\mathrm{ch}}^{2}$ (c) from the standard method (black solid markers), two-subevent method (red solid markers) and their differences (open markers). The error bars are statistical uncertainties. The decrease of $\left\langle v_{2}^{2}\right\rangle$-difference towards central collisions is attributed to two track reconstruction effects in the standard method, which does not affect the other two observables. The dashed lines indicate the upper limit of the non-flow contribution, assuming $N_{\mathrm{ch}}$-scaling of Eq. (10).
Fig. 6: Summary of the systematic uncertainties. The centrality dependence of ratios $R_{v_{2}^{2}}$ (a), $R_{(\delta p_{\mathrm{T}})^{2}}$ (b), and $R_{v_{2}^{2}\delta p_{\mathrm{T}}}$ (c) from the two-subevent method with the systematic uncertainties arising from the non-flow estimate (red bars) and other sources (black boxes).
Figure 5 also provides an estimate of non-flow based on the $N_{\mathrm{ch}}$-scaling method. We assume that the entire signals in the 80–100% centrality in two-subevent are non-flow, and then use Eq. (10) to estimate the fraction of non-flow as a function of centrality. As mentioned earlier, we employ this approach for $\left\langle v_{2}^{2}\right\rangle$ and $\left\langle v_{2}^{2}\delta p_{\mathrm{T}}\right\rangle$, where the medium effects may redistribute non-flow correlations in azimuthal angle, instead of suppressing them. This approach is unsuitable for $\left\langle(\delta p_{\mathrm{T}})^{2}\right\rangle$, for which the medium effects should always suppress the non-flow contribution. In the 0-5% most central collisions, the estimated non-flow is about 6% for $\left\langle v_{2}^{2}\right\rangle$ and only about 1.4% for $\left\langle v_{2}^{2}\delta p_{\mathrm{T}}\right\rangle$. These differences, when propagated to the ratios, are reduced for $R_{v_{2}^{2}}$, which is positive, and increased for $R_{v_{2}^{2}\delta p_{\mathrm{T}}}$, which is negative. They amount to about 2.8% for $R_{v_{2}^{2}}$ and 2.5% for $R_{v_{2}^{2}\delta p_{\mathrm{T}}}$, respectively.
The non-flow systematic uncertainties are taken as the larger of the two approaches for $R_{v_{2}^{2}}$ and $R_{v_{2}^{2}\delta p_{\mathrm{T}}}$, while for $R_{(\delta p_{\mathrm{T}})^{2}}$, the difference between standard and subevent methods is used. The total non-flow uncertainties in the 0–5% centrality are 2.8%, 3.5% and 11% for $R_{v_{2}^{2}}$, $R_{(\delta p_{\mathrm{T}})^{2}}$ and $R_{v_{2}^{2}\delta p_{\mathrm{T}}}$, respectively.
Figure 6 contrasts the non-flow systematic uncertainties with other sources of uncertainties (next section) in this analysis. In the 0–5% centrality, the non-flow uncertainties are comparable or somewhat larger than other sources, while they are subdominant in other centrality ranges.
In the literature, non-flow contributions are sometimes estimated using the HIJING model Gyulassy and Wang (1994), which has only non-flow correlations. The latter were found to follow very closely Eq. (10Bhatta et al. (2022); Wang (2024). In our second approach, instead of relying on the HIJING model, we assume this $N_{\mathrm{ch}}$-scaling behavior but use real peripheral data as the baseline for non-flow contributions. Our findings indicate that the HIJING model fails to quantitatively capture the features of non-flow. Specifically, HIJING predicts a much weaker $\Delta\eta$ dependence for $\left\langle v_{2}^{2}\right\rangle$, with only a 13% difference between the standard and two-subevent method, whereas the data indicate a 50% decrease STAR Collaboration (2023) (Fig. 25 in Ref. STAR Collaboration (2023) for $p+p$ collisions). Furthermore, we found that the values of $\left\langle v_{2}^{2}\delta p_{\mathrm{T}}\right\rangle$ predicted by HIJING are three times larger than the data in peripheral Au+Au collisions. Therefore, the non-flow estimation based on the HIJING model in Ref. Wang (2024) appears to be exaggerated. A more recent estimate Jia et al. (2024), based on a transport model incorporating full medium dynamics and Eq. (10), yields a non-flow fraction consistent with STAR data. This study also suggests a potential increase of the $R_{v_{2}^{2}}$ with $\Delta\eta$ associated with flow decorrelation effects.
Understanding non-flow correlations as a physical process has always been a work in progress. As our knowledge deepens, the non-flow uncertainties are expected to reduce. Rather than merely contributing to experimental uncertainties or even being corrected for in the data, non-flow physics should ultimately be incorporated into hydrodynamic models. Currently, these models include non-flow effects from resonance decays but lack contributions from jet fragmentation.
.9 Systematic uncertainties
@@
Systematic uncertainties include an estimate of the non-flow contributions discussed above and other sources accounting for detector effects and analysis procedure. These other sources are estimated by varying the track quality selections, the $z_{\mathrm{vtx}}$ cuts, examining the influence of pileup, comparing results from periods with different detector conditions, and closure test. The influence of track selection criteria is studied by varying the number of fit hits on the track from a minimum of 16 to 19 and by varying DCA cut from $<3$ cm to $<2.5$ cm, resulting in variations of 1–5% for $\left\langle(\delta p_{\mathrm{T}})^{2}\right\rangle$. The impacts on $\left\langle v_{2}^{2}\right\rangle$ and $\left\langle v_{2}^{2}\delta p_{\mathrm{T}}\right\rangle$ are up to $2.5\%$ and $4\%$, respectively.
@@
The influence of track reconstruction on the collision vertex is examined by comparing the results for different $|z_{\mathrm{vtx}}|$ cuts, with variations found to be $0.5$–3% for all observables. Comparisons between data-taking periods, particularly normal and reverse magnetic field runs in Au+Au collisions, show consistency within their statistical uncertainties. The influence of pileup and background events is studied by varying the cut on the correlation between $N_{\mathrm{ch}}^{\mathrm{rec}}$ and the number of hits in the TOF. The influence is found to be 1–3% for $\left\langle v_{2}^{2}\right\rangle$ and $\left\langle(\delta p_{\mathrm{T}})^{2}\right\rangle$, and reaches 2–10% for $\left\langle v_{2}^{2}\delta p_{\mathrm{T}}\right\rangle$. Comparisons are also made between the 2010 and 2011 Au+Au datasets, which have different active acceptances in the TPC. The results are largely consistent with the quoted uncertainties, although some differences are observed, particularly in the central region, where variations reach 5–10% for $\left\langle v_{2}^{2}\delta p_{\mathrm{T}}\right\rangle$.
@@
A closure test was conducted, wherein the reconstruction efficiency and its variations in $\eta$ and $\phi$ from the data were utilized to retain a fraction of the particles generated from a multi-phase transport model Lin et al. (2005). Subsequently, a track-by-track weight, as described in Eq. (9), was applied to the accepted particles. All observables are calculated using the accepted particles and compared with those obtained using the original particles. This procedure allowed us to recover $\left\langle v_{2}^{2}\right\rangle$ and $\left\langle(\delta p_{\mathrm{T}})^{2}\right\rangle$ within their statistical uncertainties. However, a 2–3% nonclosure was observed in $\left\langle v_{2}^{2}\delta p_{\mathrm{T}}\right\rangle$. Nevertheless, it’s important to note that such non-closures largely cancel when considering the ratios between U+U and Au+Au collisions.
@@
Several additional cross-checks were carried out. The track reconstruction efficiency has an approximate 5% uncertainty due to its reliance on particle type and occupancy dependence. We repeated the analysis by varying this efficiency, and the variations in the results were either less than 1% or consistent within their statistical uncertainties. The reconstructed $p_{\mathrm{T}}$ can differ from the true value due to finite momentum resolution. This effect was investigated by smearing the reconstructed $p_{\mathrm{T}}$ according to the known resolution, calculating the observable, and comparing the results with the original ones. A discrepancy of approximately 0.5% was observed for $\left\langle(\delta p_{\mathrm{T}})^{2}\right\rangle$, while other observables remained consistent within their statistical uncertainties. These effects cancel in the ratios between U+U and Au+Au collisions.